# Second- and third-order properties of multidimensional Langevin equations

### v9

@@ 6.9. Finite time interval “cubic variations”
We now briefly address the issue of measuring “cubic variations” at finite time intervals $\tau$. As mentioned before, this vanishes faster than $\tau$. We can solve a differential equation to obtain:
@@
and similarly for $j$ and $k$. We see that if second-order angular momenta vanish, then the “cubic variation” at finite time intervals $\tau$ vanishes, with corrections $a\mathcal{O}((a,b)^{2})\tau^{2}$ and $\mathcal{O}(\tau^{3})$. We may evaluate quantitative significance by treating $L$ and $b$ as uncorrelated, because they have opposite behavior under time reversal.
@@ 6.10. Comparison of experimental data and theoretical predictions
We consider the sum and difference of these quantities at $\tau=0$. The former is equal to $(\Delta t)L(x^{i}x^{j},y)$, while the latter is equal to $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. We can calculate the “ensemble covariance” of these quantities using the aforementioned prescriptions. We may consider using them for comparison at arbitrary $\tau$. However, we should not compare the sum and difference, but rather combinations of them that result in the above quantities. For example, even if $L(x^{i}x^{j},y)=0$, time-reversibility need not hold for the above quantities at finite $\tau$. Again, we may treat $L$ and $b$ as independent in the ensemble. As in the linear case, and as is analogously true for all quantities, if the theoretical values at $\tau=0$ exceed the prescriptions of the “ensemble covariance” (for either second- or third-order quantities), we may need to increase the tolerance for comparison accordingly.
@@
Again we consider the sum and difference at $\tau=0$ and generalize to arbitrary $\tau$. These are $2(\Delta t)L(x^{i}x^{j},y)$ and $4(\Delta t)b^{iy}_{k}C^{jk}$, respectively. As in the previous case, we use combinations of the sum and difference that result in the above quantities.
@@ 7.1. Linear Gaussian systems
(209) $\langle\boldsymbol{\xi}^{\pm}(t)\rangle=\mathbf{0},\quad\langle\boldsymbol{\xi}^{\pm}(t){\boldsymbol{\xi}^{\pm}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{\pm\pm}\delta(t-t^{\prime}),\quad\langle\boldsymbol{\xi}^{+}(t){\boldsymbol{\xi}^{-}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{+-}\delta(t-t^{\prime}).$
To have time-reversal symmetry, we must have $\mathbf{D}^{+-}=\mathbf{0}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=\mathbf{0}$. These imply:
@@
(211) $\langle\dot{\mathbf{x}}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=-\langle\mathbf{x}^{+}{{}\dot{\mathbf{x}}^{-}}^{\mathsf{T}}\rangle.$
We must also have the relations from before:
@@
Next, we address the time-reversibility of the second moments. It turns out that this depends only on the linear dynamics and not on the homogeneity of the diffusion matrix.
@@
(221) $\langle\xi^{y}(t)\rangle=0,\quad\langle\boldsymbol{\xi}^{\pm}(t)\xi^{y}(t^{\prime})\rangle=2\mathbf{D}^{y\pm}\delta(t-t^{\prime}),\quad\langle\xi^{y}(t)\xi^{y}(t^{\prime})\rangle=2D^{yy}\delta(t-t^{\prime}).$
For time-reversibility, we need $\mathbf{D}^{y-}=\mathbf{0}$. We write $\Delta y(\tau):=y(\tau)-y(0)$, and:
(222) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \Delta y\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&\mathbf{0}\\ \mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&\mathbf{0}\\ \mathbf{A}^{y}_{+}&\mathbf{A}^{y}_{-}&0\end{pmatrix}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \Delta y\end{pmatrix}$
Also:
(223) $\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\end{pmatrix}=\begin{pmatrix}0&\mathbf{0}^{\mathsf{T}}&\mathbf{0}^{\mathsf{T}}\\ \widehat{\mathbf{A}}^{y(1)+}_{(0)}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}\\ \widehat{\mathbf{A}}^{y(1)-}_{(0)}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}\end{pmatrix}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\end{pmatrix}$
where we wrote $\widehat{\mathbf{A}}$ for the matrix on the r.h.s. and labeled $\mathbf{x}^{\pm}\Delta y$ as $y^{(1)\pm}$. The elements are:
@@ 7.2. Third-order quantities
Here we discuss the conditions for detailed balance in the Fokker–Planck equation Eq. (86) in the presence of odd variables. Each variable $x^{i}$ has a parity $\epsilon^{i}$, and the product $\boldsymbol{\epsilon}\mathbf{x}$ is defined as $(\boldsymbol{\epsilon}\mathbf{x})^{i}=\epsilon^{i}x^{i}$ ($i$ not summed over), following [3]. In terms of the stationary distribution $p(\mathbf{x})$, the conditions for detailed balance are [3]:
@@ 7.3. Inhomogeneous diffusion
We again assume, as throughout this section, that $\mathbf{D}^{+-}=\mathbf{0}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=\mathbf{0}$. To have time-reversibility of the covariance functions, we first need that:
@@
(245) $\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix}=\begin{pmatrix}0&&&&\\ {\widehat{\mathbf{A}}}^{y(1)+}_{(0)}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(1)-}_{(0)}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(2)+}_{(0)}&\widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}\\ {\widehat{\mathbf{A}}}^{y(2)-}_{(0)}&\widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}\end{pmatrix}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix}.$
To have time-reversibility of the third-order covariance functions, we must have for $n\geq 1$:
@@ 7.4. Nonlinear drift
which can be proven by induction on $n$. We also have, by orthogonality of the Hermite polynomials and symmetry:
@@
(251) $\langle\mathbf{x}^{(4)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ $=\mathcal{O}(a,b),$
which implies, for $n\geq 1$:
(252) ${{}\widehat{\mathbf{A}}^{n}}^{(3)}$ $=\mathcal{O}((a,b)^{2})$
(253) ${{}\widehat{\mathbf{A}}^{n}}^{(4)}$ $=\mathcal{O}(a,b).$
@@
Again, the base case $n=1$ for Eq. (255) follows from Eqs. (236)–(237). To carry out the proof by induction for Eq. (254)–(255) assuming the base case $n=1$, we need to use, for $i\geq 1$:
(256) ${{}\widehat{\mathbf{A}}^{n}}^{(i+1)}_{(i)}=\mathcal{O}(a,b)$
which allows us to write:
(257) ${{}\widehat{\mathbf{A}}^{n}}^{(i)}_{(i)}={{{}\widehat{\mathbf{A}}}^{(i)}_{(i)}}^{n}+a\mathcal{O}(a,b).$
We also need to use that $\widehat{\mathbf{A}}^{(2)}_{(2)}=\mathbf{A}\otimes\mathbf{A}$ which means that the same rules of time-reversibility that apply to $\mathbf{x}$ and $\mathbf{A}$ also apply to $\mathbf{x}^{(2)}$ and $\widehat{\mathbf{A}}^{(2)}_{(2)}$ (respectively).
@@ 7.5. Integrated variables
For an integrated variable $y$, we have for $n\geq 1$, $i\geq 1$, $j\geq 1$:
(258) ${{}\widehat{\mathbf{A}}^{n}}^{y(i-1)}_{(i+j)}=\mathcal{O}(a^{j})[1+a\mathcal{O}(a,b)].$
Also, for $n\geq 1$, $i\geq 0$, $j\geq 0$:
@@
We also have:
(261) ${{}\widehat{\mathbf{A}}^{n}}^{(1)}_{(0)}$ $=a^{2}\mathcal{O}(a,b),$
(262) ${{}\widehat{\mathbf{A}}^{n}}^{y(0)}_{(0)}$ $=a^{2}\mathcal{O}(a,b).$
From orthogonality of Hermite polynomials and symmetry, we have:
(263) $\langle\Delta y(\tau)\mathbf{x}^{(2)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ $=\mathcal{O}((a,b)^{2}),$
@@
(265) ${{}\widehat{\mathbf{A}}^{n}}^{y(2)}$ $=\mathcal{O}((a,b)^{2}),$
(266) ${{}\widehat{\mathbf{A}}^{n}}^{y(3)}$ $=\mathcal{O}(a,b).$
@@
(272) ${{}\widehat{\mathbf{A}}^{n}}^{(2)}_{(0)}$ $=a\mathcal{O}(a,b),$
(273) $\left|{{}\widehat{\mathbf{A}}^{n}}^{(m)}_{(0)}\right|$ $\leq\mathcal{O}((a,b)^{2}),\quad m\geq 4$
@@
(276) ${{}\widehat{\mathbf{A}}}^{(3)}_{(0)}+{{}\widehat{\mathbf{A}}}^{(3)}_{(3)}\langle\mathbf{x}^{(3)}\rangle=a\mathcal{O}((a,b)^{2}).$

### v11

@@ 6.9. Finite time interval “cubic variations”
We now briefly address the issue of measuring “cubic variations” at finite time intervals $\tau$. As mentioned before, this vanishes faster than $\mathcal{O}(\tau)$. We can solve a differential equation to obtain:
@@
and similarly for $j$ and $k$. We see that if second-order angular momenta vanish, then the “cubic variation” at finite time intervals $\tau$ vanishes, with corrections $a\mathcal{O}((a,b)^{2})\tau^{2}$ and $\mathcal{O}(\tau^{3})$. We may evaluate quantitative significance by treating $L$ and $b$ as uncorrelated for the “ensemble covariance”, because they acquire opposite signs under time reversal.
@@ 6.10. Comparison of experimental data and theoretical predictions
We consider the sum and difference of these quantities at $\tau=0$. The former is equal to $(\Delta t)L(x^{i}x^{j},y)$, while the latter is equal to $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. We can calculate the “ensemble covariances” of these quantities using the aforementioned prescriptions. We may consider using them for comparison at arbitrary $\tau$. However, we may be interested in comparing the above quantities directly. They may be obtained by taking one-half the sum and difference of the sum and difference. We may treat the sum and difference of the above quantities as uncorrelated in the “ensemble covariance” as they acquire opposite signs under time reversal. Thus we conclude that the above quantities should be compared according to one-fourth the sum of the “ensemble covariances” of $(\Delta t)L(x^{i}x^{j},y)$ and $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. As in the linear case, and as is analogously true for all quantities, if the theoretical values at $\tau=0$ exceed the prescriptions of the “ensemble variances” (for either second- or third-order quantities), we may need to increase the tolerance for comparison accordingly.
@@
Again we consider the sum and difference at $\tau=0$ and generalize to arbitrary $\tau$. These are $2(\Delta t)L(x^{i}x^{j},y)$ and $4(\Delta t)b^{iy}_{k}C^{jk}$, respectively. As in the previous case, we may use combinations of the sum and difference that result in the above quantities.
@@ 7.1. Linear Gaussian systems
(218) $\langle\boldsymbol{\xi}^{\pm}(t)\rangle=\mathbf{0},\quad\langle\boldsymbol{\xi}^{\pm}(t){\boldsymbol{\xi}^{\pm}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{\pm\pm}\delta(t-t^{\prime}),\quad\langle\boldsymbol{\xi}^{+}(t){\boldsymbol{\xi}^{-}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{+-}\delta(t-t^{\prime}),$
where $\boldsymbol{\xi}^{\pm}$ are Gaussian random noise and all coefficients are constant. To have time-reversal symmetry, we must have $\mathbf{D}^{+-}=\mathbf{0}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=\mathbf{0}$. These imply:
@@
(220) $\langle\dot{\mathbf{x}}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=-\langle\mathbf{x}^{+}{{}\dot{\mathbf{x}}^{-}}^{\mathsf{T}}\rangle$
(with time-derivatives interpreted in Itô sense). We must also have the relations from before:
@@
Next, we address the time-reversibility of the second moments. Similarly as before, this depends only on the linear dynamics and not on the homogeneity of the diffusion matrix.
@@
(230) $\langle\xi^{y}(t)\rangle=0,\quad\langle\boldsymbol{\xi}^{\pm}(t)\xi^{y}(t^{\prime})\rangle=2\mathbf{D}^{y\pm}\delta(t-t^{\prime}),\quad\langle\xi^{y}(t)\xi^{y}(t^{\prime})\rangle=2D^{yy}\delta(t-t^{\prime}),$
where $\xi^{y}$ is Gaussian random noise and all coefficients are constant. For time-reversibility, we need $\mathbf{D}^{y-}=\mathbf{0}$. We write $\Delta y(\tau):=y(\tau)-y(0)$, and:
(231) $\left\langle{\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \Delta y\end{pmatrix}\mid\mathbf{x}}\right\rangle=\begin{pmatrix}\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&\mathbf{0}\\ \mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&\mathbf{0}\\ \mathbf{A}^{y}_{+}&\mathbf{A}^{y}_{-}&0\end{pmatrix}\begin{pmatrix}\mathbf{x}^{+}\\ \mathbf{x}^{-}\\ \Delta y\end{pmatrix}.$
Writing $\widehat{\mathbf{A}}$ for the matrix appearing on the r.h.s., the conditional expectation of the vector appearing in the above differential equation given $\mathbf{x}(0)$ is equal to $e^{\widehat{\mathbf{A}}\tau}$ multiplied by its value at $\tau=0$. Also:
(232) $\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\end{pmatrix}=\begin{pmatrix}0&\mathbf{0}&\mathbf{0}\\ \widehat{\mathbf{A}}^{y(1)+}_{(0)}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}\\ \widehat{\mathbf{A}}^{y(1)-}_{(0)}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}\end{pmatrix}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\end{pmatrix}$
where we labeled $\mathbf{x}^{\pm}\Delta y$ as $y^{(1)\pm}$. Again, we write $\widehat{\mathbf{A}}$ for the matrix on the r.h.s. and the vector [^15] A collection of quantities, not a true physical vector obeying the transformation laws under a change of basis. appearing in the above differential equation is equal to $e^{\widehat{\mathbf{A}}\tau}$ multiplied by its value when $\tau=0$. The elements are:
@@ 7.2. Third-order quantities
Here we discuss the conditions for detailed balance in the Fokker–Planck equation Eq. (93) in the presence of odd variables. Each variable $x^{i}$ has a parity $\epsilon^{i}=\pm 1$, where under time-reversal $x^{i}\to\epsilon^{i}x^{i}$ ($i$ not summed over), and the product $\boldsymbol{\epsilon}\mathbf{x}$ is defined as $(\boldsymbol{\epsilon}\mathbf{x})^{i}=\epsilon^{i}x^{i}$ ($i$ not summed over), following [3]. In terms of the stationary distribution $p(\mathbf{x})$, the conditions for detailed balance are [3]:
@@ 7.3. Inhomogeneous diffusion
We again assume, as throughout this section, that $\mathbf{D}^{+-}=\mathbf{0}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=\mathbf{0}$, and we denote the matrix on the r.h.s. by $\widehat{\mathbf{A}}$. To have time-reversibility of the covariance functions, we first need that:
@@
(254) $\frac{\mathrm{d}}{\mathrm{d}\tau}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix}=\begin{pmatrix}0&&&&\\ {\widehat{\mathbf{A}}}^{y(1)+}_{(0)}&\mathbf{A}^{+}_{+}&\mathbf{A}^{+}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(1)-}_{(0)}&\mathbf{A}^{-}_{+}&\mathbf{A}^{-}_{-}&&\\ {\widehat{\mathbf{A}}}^{y(2)+}_{(0)}&\widehat{\mathbf{A}}^{(2)+}_{+}&\widehat{\mathbf{A}}^{(2)+}_{-}&\widehat{\mathbf{A}}^{(2)+}_{(2)+}&\widehat{\mathbf{A}}^{(2)+}_{(2)-}\\ {\widehat{\mathbf{A}}}^{y(2)-}_{(0)}&\widehat{\mathbf{A}}^{(2)-}_{+}&\widehat{\mathbf{A}}^{(2)-}_{-}&\widehat{\mathbf{A}}^{(2)-}_{(2)+}&\widehat{\mathbf{A}}^{(2)-}_{(2)-}\end{pmatrix}\begin{pmatrix}1\\ \langle\mathbf{x}^{+}\Delta y\rangle\\ \langle\mathbf{x}^{-}\Delta y\rangle\\ \langle\mathbf{x}^{(2)+}\Delta y\rangle\\ \langle\mathbf{x}^{(2)-}\Delta y\rangle\end{pmatrix},$
where the superscript $y(2)\pm$ stands for $\mathbf{x}^{(2)\pm}\Delta y$. To have time-reversibility of the third-order covariance functions, we must have for $n\geq 1$:
@@ 7.4. Nonlinear drift
(259) ${{}\widehat{\mathbf{A}}^{n}}^{(i+1)}_{(i)}$ $=\mathcal{O}(a,b),$
which can be proven by induction on $n$ [^16] The index “(1)” refers to $\mathbf{x}$ and is sometimes omitted, and the index “(0)” refers to the constant 1, so that $\langle\mathrm{d}\mathbf{x}^{(i)}(\tau)/\mathrm{d}\tau\mid\mathbf{x}(\tau)\rangle=\sum_{j=0}^{\infty}{\widehat{\mathbf{A}}}^{(i)}_{(j)}\mathbf{x}^{(j)}(\tau)$ and $\langle\mathbf{x}^{(i)}(\tau)\mid\mathbf{x}(0)\rangle=\sum_{j=0}^{\infty}(e^{\widehat{\mathbf{A}}\tau})^{(i)}_{(j)}\mathbf{x}^{(j)}(0)$.. Also:
(260) $\langle\mathbf{x}(\tau)\rangle$ $=a^{2}\mathcal{O}(a,b)$
(261) $\langle\mathbf{x}^{(2)}(\tau)\rangle$ $=a\mathcal{O}(a,b)$
@@
(263) $\langle\mathbf{x}^{(4)}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ $=\mathcal{O}(a,b).$
Now, we have:
(264) $\langle\mathbf{x}(\tau)\rangle=(e^{\widehat{\mathbf{A}}\tau})^{(1)}_{(0)}+(e^{\widehat{\mathbf{A}}\tau})^{(1)}_{(1)}\langle\mathbf{x}(0)\rangle+(e^{\widehat{\mathbf{A}}\tau})^{(1)}_{(2)}\langle\mathbf{x}^{(2)}(0)\rangle+(e^{\widehat{\mathbf{A}}\tau})^{(1)}_{(3)}\langle\mathbf{x}^{(3)}(0)\rangle+\cdots,$
with:
(265) $\langle\mathbf{x}(0)\rangle=a^{2}\mathcal{O}(a,b),\quad\langle\mathbf{x}^{(2)}(0)\rangle=a\mathcal{O}(a,b),\quad\langle\mathbf{x}^{(3)}(0)\rangle=\mathcal{O}(a,b).$
We Taylor expand $e^{\widehat{\mathbf{A}}\tau}$ and use Eqs. (258)–(259) to conclude that for $n\geq 1$:
(266) ${{}\widehat{\mathbf{A}}^{n}}^{(1)}_{(0)}=a^{2}\mathcal{O}(a,b).$
Similarly, we have:
(267) ${{}\widehat{\mathbf{A}}^{n}}^{(2)}_{(0)}$ $=a\mathcal{O}(a,b)$
(268) ${{}\widehat{\mathbf{A}}^{n}}^{(3)}_{(1)}$ $=\mathcal{O}((a,b)^{2})$
(269) ${{}\widehat{\mathbf{A}}^{n}}^{(4)}_{(1)}$ $=\mathcal{O}(a,b).$
@@
Again, the base case $n=1$ for Eq. (271) follows from Eqs. (245)–(246). To carry out the proof by induction for Eq. (270)–(271) assuming the base case $n=1$, we need to use, for $i\geq 1$ and $n\geq 1$:
(272) ${{}\widehat{\mathbf{A}}^{n}}^{(i)}_{(i)}={{{}\widehat{\mathbf{A}}}^{(i)}_{(i)}}^{n}+a\mathcal{O}(a,b)$
which follows from Eqs. (258)–(259). We also need to use that $\widehat{\mathbf{A}}^{(2)}_{(2)}=\mathbf{A}\otimes\mathbf{A}$ which means that the same rules of time-reversibility that apply to $\mathbf{x}$ and $\mathbf{A}$ also apply to $\mathbf{x}^{(2)}$ and $\widehat{\mathbf{A}}^{(2)}_{(2)}$ (respectively).
@@ 7.5. Integrated variables
For an integrated variable $y$ obeying Eq. (173), we have for $n\geq 1$, $i\geq 1$, $j\geq 1$:
(273) ${{}\widehat{\mathbf{A}}^{n}}^{y(i-1)}_{(i+j)}=\mathcal{O}(a^{j})[1+a\mathcal{O}(a,b)],$
where the superscript “$y(i)$” refers to $\mathbf{x}^{(i)}\Delta y$ (“$y(0)$” being synonymous with “$y$”). Also, for $n\geq 1$, $i\geq 0$, $j\geq 0$:
@@
We also have $\langle\dot{y}\rangle=a^{2}\mathcal{O}(a,b)$, which implies that $\langle\Delta y(\tau)\rangle=a^{2}\mathcal{O}(a,b)$, and hence:
(276) ${{}\widehat{\mathbf{A}}^{n}}^{y}_{(0)}=a^{2}\mathcal{O}(a,b).$
Additionally:
(277) $\langle\Delta y(\tau)\mathbf{x}^{(2)}(\tau)\rangle$ $=\mathcal{O}(a,b)$
@@
(279) ${{}\widehat{\mathbf{A}}^{n}}^{y(2)}_{(0)}$ $=\mathcal{O}(a,b),$
(280) ${{}\widehat{\mathbf{A}}^{n}}^{y(3)}_{(1)}$ $=\mathcal{O}(a,b).$
@@
(286) ${{}\widehat{\mathbf{A}}^{n}}^{(4)}_{(0)}$ $=\mathcal{O}((a,b)^{2}),$
(287) ${{}\widehat{\mathbf{A}}^{n}}^{(5)}_{(0)}$ $=\mathcal{O}(a,b),$
@@
(290) $0=\frac{\mathrm{d}}{\mathrm{d}\tau}\langle\mathbf{x}^{(3)}\rangle={{}\widehat{\mathbf{A}}}^{(3)}_{(0)}+{{}\widehat{\mathbf{A}}}^{(3)}_{(3)}\langle\mathbf{x}^{(3)}\rangle+a\mathcal{O}((a,b)^{2}).$